Determine values of n for which Dihedral group Dn is nilpotent.


1
Expert's answer
2021-02-03T01:18:57-0500

We denote by "D_{2n}" the dihedral group of order "2n".

"D_{2n}" is nilpotent if n a power of 2: "n=2^m" with "m\geq0". In fact,

Suppose "D_{2 n}" is nilpotent. Let "p" be prime number which divides "n" . Then "r^{n \/ p}" is an element of "D_{2 n}" order "p", so "r^{n / p} \neq r^{-n / p}" . Let "|s|=2" and "|r^{n \/ p}|=p" are relatively prime, so that, "s r^{n \/ p}=r^{n \/ p} s" ; a contradiction. Thus no primes divide "n" , and we have "n=2^{k}".


reciprocally, we proceed by induction on "k" where "n=2^{k}".

For "k=0", "D_{2.2^{0}} \cong Z_{2}" is abelian, hence nilpotent.

Suppose "D_{2.2^{k}}" is nilpotent. We have "Z\left(D_{2 \cdot 2^{k+1}}\right)=\left\langle r^{2^{k}}\right\rangle," and so, "D_{2 \cdot 2^{k+1}} / Z\left(D_{2 \cdot 2^{k+1}}\right) \cong D_{2 \cdot 2^{k}}" is nilpotent. Thus, "D_{2 \cdot 2^{k+1}}" is nilpotent.